Use Newton's method to find all roots of the equation correct to six decimal places. (Enter your answers as a comma-separated list.)

x4 = 3 + x

x =

Find f.

f ''(x) = 4 − 12x,    f(0) = 6,    f(2) = 10

f(x)

Find the limit, if it exists. (If an answer does not exist, enter DNE.)

lim x→−∞

x +

x2 + 2x
Answer #1

= 1.5 2 Jb1a) = 2*?- 2x + en +(2 Sol : 24- = 3+x Let f(x)=x²-3-4 >$()= 4*3-1 By Newtons method! Matt =an finn) Now f(t) =-Soln 3 : lim (a tu ²+2x) N-00 ntattan = lim x(1+x+2) . n-bo a(3+1) = lim M- C Am: CScanned with CamScanner